A train travelingat 50 mph leaves for a certain town. After 2 hours, a van traveling at 60 mph leaves for the same town and arrives at the same time as the train. If both the train and the van traveled in a straight line, how far is the town from where they started?
Answer by lwsshak3(6519)

A train travelingat 50 mph leaves for a certain town. After 2 hours, a van traveling at 60 mph leaves for the same town and arrives at the same time as the train. If both the train and the van traveled in a straight line, how far is the town from where they started?
..
let x=time of travel for train
x-2 =time of travel for van
distance=speed*time of travel (same for train and van)
50x=60(x-2
50x=60x-120
10x=120
x=12 hours
distance=50*12=60(12-2)=600 miles
ans:
The town is 600 miles from where they started.